Determine the required concentration (in percent by mass) for an aqueous ethylene glycol (C2H6O2) solution to have a boiling point of 106.2 ∘C
1
Expert's answer
2020-03-12T07:44:04-0400

"\Delta"Tboil = i⋅Kb⋅Cm

Cm = ΔTboil / Kb = 6.2 / 0.512 = 12.11 (mol/kgH2O)

mC2H6O2 = nC2H6O2⋅MC2H6O2 = 12.11⋅62 = 750.82 (g)

msolution = 1000 + 750.82 = 1750.82 (g)

"\omega"C2H6O2 = 100⋅mC2H6O2 / msolution = 100⋅750.82 / 1750.82 = 42.88 %

Answer: 42.88 %.